DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, October 20, 2004 Padelford C-36 3:50 pm

Spectrum and pseudospectrum for non-selfadjoint operators

Johannes Sjöstrand (L’Ecole Polytechnique, France)

We describe some recent developments for non-selfadjoint problems: Spectral asymptotics in dimension 1 and 2, pseudospectrum (resolvent growth), spectral instability and evolution problems including the Fokker-Planck operator.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.